OpenStudy (anonymous):

What are the total possible outcomes of rolling a pair of dice? Note that the pair 3,4 is not the same as the pair 4,3. What is the probability of rolling a 7 with a pair of dice? P(7) = No. of successful outcomes/Total No. of outcomes

OpenStudy (sidsiddhartha):

there are 36 possible outcomes when ure rolling a pair of dice now 7 could be achieved by =(4,3),(3,4),(5,2),(2,5),(6,1),(1,6) so in total 6 ways now use the formula

OpenStudy (sidsiddhartha):

do u get it? @ksmith0974

OpenStudy (anonymous):

working it thx..I'll let you know

OpenStudy (sidsiddhartha):

so what's ur answer?

OpenStudy (sidsiddhartha):

no of successful outcomes=6 and no of possible outcomes=36

OpenStudy (sidsiddhartha):

so what is P(7) can u do it?

OpenStudy (anonymous):

Probability of 7 = 6/36 = 1/6

OpenStudy (sidsiddhartha):

yeah

OpenStudy (anonymous):

thx for the help

OpenStudy (sidsiddhartha):

^_^